Calculate the pH of a solution prepared by mixing equal volumes of 0.18M methylamine (CH3NH2, Kb = 3.7
Answer #1

                                         Pkb     = -log3.7*10-7

                                                   = 3.37

                                       

                                        POH      = PKb +log[salt]

                                                                     [Base]

                                                    = 3.37+ LOG(0.65/0.18

                                                   = 3.37+0.557 = 3.927

                                        PH      = 14- POH  

                                                     = 14-3.92

                                                    = 10.07